# Thermodynamic topological classes of the rotating, accelerating black holes

### v1

@@ 3.2 Kerr-AdS C-metric black hole
It is easy to verify that the above thermodynamic quantities obey the differential first law and integral Bekenstein-Smarr mass formula simultaneously,
@@
In order to obtain the Gibbs free energy of the Kerr-AdS C-metric black hole, one can calculate the Euclidean action integral JHEP0517116
$\mathcal{I}_{E}=\frac{1}{16\pi}\int_{M}d^{4}x\sqrt{g}\left(R+\frac{6}{l^{2}}\right)+\frac{1}{8\pi}\int_{\partial M}d^{3}x\sqrt{h}\left[\mathcal{K}-\frac{2}{l}-\frac{l}{2}\mathcal{R}(h)\right]\,,$ (26)
where the extrinsic curvature $\mathcal{K}$ and the Ricci scalar $\mathcal{R}(h)$ correspond to the boundary metric $h_{\mu\nu}$. To solve the divergence, the action includes not only the Einstein-Hilbert term but also the Gibbons-Hawking boundary term as well as the corresponding AdS boundary counterterms PRD60-104001 ; PRD60-104026 ; PRD60-104047 ; CMP208-413 ; CMP217-595 .
Using the results provided in Eq. (LABEL:ThermKerrAdSC) and the expression of the mass parameter
$m=\frac{r_{h}^{2}+a^{2}}{2r_{h}}-\frac{r_{h}(r_{h}^{2}+a^{2})}{2l^{2}(A^{2}r_{h}^{2}-1)}\,,$
the Gibbs free energy for the Kerr-AdS C-metric black hole is given by
$G=\frac{\mathcal{I}_{E}}{\beta}=\frac{m(1-A^{2}a^{2}-2A^{2}l^{2}\Xi)}{2\alpha K(1+A^{2}a^{2})}-\frac{r_{h}(r_{h}^{2}+a^{2})}{2\alpha Kl^{2}(1-A^{2}r_{h}^{2})^{2}}=\frac{M}{2}-VP=M-TS-\Omega\,J\,.$ (27)
It is easy to see that similar to the case of the Kerr C-metric black hole in the previous subsection, the conical singularity also has no effect on the calculation of the Gibbs free energy.
To determine the thermodynamic topological number of the Kerr-AdS C-metric black hole, we must first obtain the expression for the generalized off-shell Helmholtz free energy. The Helmholtz free energy is defined as
$F=G+\Omega\,J=M-TS\,.$ (28)
@@
Figure 4: Zero points of the vector $\phi^{r_{h}}$ shown in the $r_{h}-\tau$ plane with $a/r_{0}=1$, $Ar_{0}=1$ and $Pr_{0}^{2}=0.5$. There is one thermodynamically stable and one thermodynamically unstable Kerr-AdS C-metric black hole branch for any value of $\tau$. Obviously, the topological number is: $W=0$.
Figure 5: The arrows represent the unit vector field $n$ on a portion of the $r_{h}-\Theta$ plane for the Kerr-AdS C-metric black hole with $a/r_{0}=1$, $Ar_{0}=1$, $Pr_{0}^{2}=0.5$ and $\tau/r_{0}=20$. The zero points (ZPs) marked with black dots are at $(r_{h}/r_{0},\Theta)=(0.33,\pi/2)$, $(2.23,\pi/2)$, for ZP1 and ZP2, respectively. The blue contours $C_{i}$ are closed loops surrounding the zero points.
@@
Furthermore, in light of another intriguing finding presented in Ref. PRD108-084041 , which demonstrate that the topological number of an non-rotating, accelerating AdS black hole is consistently lower by $-1$ compared to its non-accelerating counterpart, a broader conjecture within the framework of general relativity is proposed: the topological number of an accelerating AdS black hole consistently differs by minus one from that of its corresponding non-accelerating AdS black hole. This conjecture will be further explored in section 4.2, where we calculate the topological number of the KN-AdS C-metric black hole and compare it to that of the KN-AdS black hole, thereby affirming the validity and relevance of our proposal.
@@
4 Topological classes of rotating charged accelerating black hole
@@
4.1 KN C-metric black hole
@@
$Q=\frac{q}{K}\,,~{}\qquad\Phi=\frac{qr_{h}}{\alpha(r_{h}^{2}+a^{2})}\,,~{}\qquad\Xi=1+A^{2}(a^{2}+q^{2})\,.$
Nevertheless, we can verify that both the differential and integral mass formulas are completely satisfied
@@
For the KN C-metric black hole, the expression of the Gibbs free energy can be calculated by the Euclidean action integral
$\mathcal{I}_{E}=\frac{1}{16\pi}\int_{M}d^{4}x\sqrt{g}\big{(}R-F^{2}\big{)}+\frac{1}{8\pi}\int_{\partial{}M}d^{3}x\sqrt{h}(\mathcal{K}-\mathcal{K}_{0})\,.$ (37)
where $h$ represents the determinant of the induced metric $h_{ij}$, while $\mathcal{K}$ denotes the extrinsic curvature of the boundary. The term $\mathcal{K}_{0}$ indicates the subtracted value associated with the massless C-metric solution, which is utilized as the reference background. Therefore, the result of the Gibbs free energy straightforwardly reads
$G=\frac{\mathcal{I}_{E}}{\beta}=\frac{m(1-A^{2}a^{2})}{2\alpha K(1+A^{2}a^{2})}-\frac{q^{2}r_{h}}{2\alpha K(r_{h}^{2}+a^{2})}=\frac{M-\Phi Q}{2}=M-TS-\Omega\,J-\Phi Q\,,$ (38)
with the thermodynamic quantities in (LABEL:ThermKNC) as required. In addition, it indicate that the conical singularity also has no effect on the calculation of the Gibbs free energy in this case.
Next, we will discuss the topological number of the KN C-metric black hole. We note that the Helmholtz free energy is given by
$F=G+\Omega\,J+\Phi Q=M-TS\,.$ (39)
It is simple to derive the generalized off-shell Helmholtz free energy as
@@
Figure 6: Zero points of the vector $\phi^{r_{h}}$ shown in the $r_{h}-\tau$ plane with $a/r_{0}=1$, $q/r_{0}=1$ and $Ar_{0}=0.5$. There is one thermodynamically stable KN C-metric black hole for any value of $\tau$. Obviously, the topological number is: $W=1$.
Figure 7: The arrows represent the unit vector field $n$ on a portion of the $r_{h}-\Theta$ plane for the KN C-metric black hole with $a/r_{0}=1$, $q/r_{0}=1$, $Ar_{0}=0.5$ and $\tau/r_{0}=20$. The zero point (ZP) marked with black dot is at $(r_{h}/r_{0},\Theta)=(3.56,\pi/2)$. The blue contours $C$ is closed loop enclosing the zero point.
Considering $a/r_{0}=1$, $q/r_{0}=1$ and $Ar_{0}=0.5$ for the KN C-metric black hole, we plot the zero points of $\phi^{r_{h}}$ in the $r_{h}-\tau$ plane in Fig. 6, and the unit vector field $n$ on a portion of the $\Theta-r_{h}$ plane with $\tau/r_{0}=20$ in Fig. 7. Evidently, for any value of $\tau$, there exists a unique thermodynamically stable KN C-metric black hole. As shown in Fig. 7, the zero point is located at $(r_{h}/r_{0},\Theta)=(3.56,\pi/2)$. By analyzing the local properties of this zero point, we can readily determine the topological number $W=1$ for the KN C-metric black hole. Comparing this result with the corresponding finding for the Kerr C-metric black hole in section 3.1, where $W=1$ as well, it becomes evident that the presence of the electric charge parameter does not influence the topological number of rotating, accelerating black holes. This observation further supports our initial conjecture, as proposed in Ref. PRD107-024024 , that the topological number remains unaffected by the charge in rotating black holes–a conclusion that not only extends to the case of rotating black holes in gauged supergravity theories PLB856-138919 , but also to the case of rotating, accelerating black holes. In addition, by comparing with the corresponding result for the KN black hole ($W=0$), we further validate the conjecture proposed in section 3.1: the difference in topological numbers between any asymptotically flat accelerating black hole and its non-accelerating counterpart consistently equals one. This conclusion also holds true in the case of rotating, accelerating charged black holes.
4.2 KN-AdS C-metric black hole
In this subsection, we consider the general KN-AdS C-metric black hole case. The metric, the Abelian gauge potential, and the metric functions are already given in Eqs. (4)-(2). The thermodynamic quantities of the KN-AdS C-metric black hole are JHEP0419096
@@ 4.2 KN-AdS C-metric black hole
We now examine the Gibbs free energy of the KN-AdS C-metric black hole through the Euclidean action integral method. The Euclidean action is expressed as
$\mathcal{I}_{E}=\frac{1}{16\pi}\int_{M}d^{4}x\sqrt{g}\Big{(}R+\frac{6}{l^{2}}-F^{2}\Big{)}+\frac{1}{8\pi}\int_{\partial M}d^{3}x\sqrt{h}\Big{[}\mathcal{K}-\frac{2}{l}-\frac{l}{2}\mathcal{R}(h)\Big{]}\,,$ (47)
where $\mathcal{K}$ and $\mathcal{R}(h)$ denote the extrinsic curvature and the Ricci scalar of the boundary metric $h_{\mu\nu}$, respectively. In addition to the standard Einstein-Hilbert term, the action includes the Gibbons-Hawking boundary term and the corresponding AdS boundary counterterms, which are introduced to remove divergences. Therefore, the Gibbs free energy is simply given by
$G$ $=$ $\frac{\mathcal{I}_{E}}{\beta}=\frac{m(1-A^{2}a^{2}-2A^{2}l^{2}\Xi)}{2\alpha K(1+A^{2}a^{2})}-\frac{r_{h}(r_{h}^{2}+a^{2})}{2\alpha Kl^{2}(1-A^{2}r_{h}^{2})^{2}}-\frac{q^{2}r_{h}}{2\alpha K(r_{h}^{2}+a^{2})}$ (48)
$=$ $\frac{M-\Phi Q}{2}-VP=M-TS-\Omega\,J-\Phi Q\,.$
where $\beta=T^{-1}$ represents the time coordinate’s interval. The final two equalities hold true when considering the thermodynamic variables from Eq. (LABEL:ThermKNAdSC). Moreover, it is evident that the conical singularity does not influence the calculation of the Gibbs free energy in this case.
Figure 8: Zero points of the vector $\phi^{r_{h}}$ shown in the $r_{h}-\tau$ plane with $a/r_{0}=1$, $q/r_{0}=1$, $Ar_{0}=1$ and $Pr_{0}^{2}=0.5$. There is one thermodynamically stable and one thermodynamically unstable KN-AdS C-metric black hole branch for any value of $\tau$. Obviously, the topological number is: $W=0$.
In the following, we will investigate the topological number of the KN-AdS C-metric black hole. The Helmholtz free energy simply reads
$F=G+\Omega\,J+\Phi Q=M-TS\,.$ (49)
Then the generalized off-shell Helmholtz free energy is given by
@@
$\phi^{r_{h}}$ $=$ $\frac{\sqrt{1+A^{2}(a^{2}+q^{2})}}{6K[1-8\pi Pa^{2}/3+A^{2}(a^{2}+q^{2})]}\sqrt{1+\frac{A^{2}[-3+8\pi Pa^{2}-3A^{2}(a^{2}+q^{2})]}{8\pi P}}\Bigg{\{}3-\frac{3q^{2}}{r_{h}^{2}}$ (51)
$-\frac{8\pi Pr_{h}^{2}(A^{2}r_{h}^{2}-3)}{(A^{2}r_{h}^{2}-1)^{2}}+a^{2}\Bigg{[}4\pi P\bigg{(}\frac{1}{(Ar_{h}-1)^{2}}+\frac{1}{(Ar_{h}+1)^{2}}\bigg{)}\Bigg{]}\Bigg{\}}-\frac{2\pi A^{2}r_{h}(r_{h}^{2}+a^{2})}{K\tau(A^{2}r_{h}^{2}-1)^{2}}$
$-\frac{2\pi r_{h}}{K\tau(1-A^{2}r_{h}^{2})}\,,$
@@
$\tau$ $=$ $-\frac{8\sqrt{2P}\pi^{\frac{3}{2}}r_{h}^{3}(1+A^{2}a^{2})\big{[}3-8\pi Pa^{2}+3A^{2}(a^{2}+q^{2})\big{]}}{\sqrt{\big{[}1+A^{2}(a^{2}+q^{2})\big{]}\big{[}8\pi P+A^{2}(8\pi Pa^{2}-3)-3A^{4}(a^{2}+q^{2})\big{]}}}\Big{\{}-3(r_{h}^{2}$ (53)
$+8\pi Pr_{h}^{4})+3q^{2}(A^{2}r_{h}^{2}-1)^{2}+A^{2}r_{h}^{4}\big{[}r_{h}^{2}(8\pi P-3A^{2})+6\big{]}+a^{2}\big{[}3(A^{2}r_{h}^{2}-1)^{2}$
$-8\pi Pr_{h}^{2}(1+A^{2}r_{h}^{2})\big{]}\Big{\}}^{-1}$
@@
Figure 9: The arrows represent the unit vector field $n$ on a portion of the $r_{h}-\Theta$ plane for the KN-AdS C-metric black hole with $a/r_{0}=1$, $q/r_{0}=1$, $Ar_{0}=1$, $Pr_{0}^{2}=0.5$ and $\tau/r_{0}=20$. The zero points (ZPs) marked with black dots are at $(r_{h}/r_{0},\Theta)=(0.42,\pi/2)$, $(2.17,\pi/2)$, for ZP1 and ZP2, respectively. The blue contours $C_{i}$ are closed loops surrounding the zero points.
In Figs. 8 and 9, taking $a/r_{0}=1$, $q/r_{0}=1$, $Ar_{0}=1$, $Pr_{0}^{2}=0.5$ for the KN-AdS C-metric black hole, we plot the zero points of $\phi^{r_{h}}$ in the $r_{h}-\tau$ plane and the unit vector field $n$ with $\tau=20r_{0}$, respectively. Note that for these values of $a/r_{0}$, $q/r_{0}$, $Ar_{0}$ and $Pr_{0}^{2}$, there are one thermodynamically stable and one thermodynamically unstable KN-AdS C-metric black hole for any value of $\tau$. In Fig. 9, one can observe that two zero points are located at $(r_{h}/r_{0},\Theta)=(0.42,\pi/2)$ and $(2.17,\pi/2)$, respectively. As a result, the topological number $W=0$ for the KN-AdS C-metric black hole can be explicitly established in Figs. 8 and 9 via the local property of the zero point, which is different from that of the four-dimensional KN-AdS black hole ($W=1$) PRD107-084002 . Therefore, we further verify that the conjecture proposed in section 3.2–”the topological number of an asymptotically AdS, accelerating black hole consistently differs by minus one from that of its corresponding asymptotically AdS, non-accelerating black hole”–is also applicable to the case of rotating, accelerating charged AdS black holes. Furthermore, since the KN-AdS C-metric black hole retains the same topological number as the KN black hole ($W=0$), we show that although both the acceleration parameter and the negative cosmological constant individually increase the topological number by one, their combined effect in the rotating charged black hole solution counteracts this increase.
5 Concluding remarks

### v2

@@ III.2 Kerr-AdS C-metric black hole
where $V$ is the thermodynamic volume. It is easy to verify that the above thermodynamic quantities obey the differential first law and integral Bekenstein-Smarr mass formula simultaneously,
@@
Figure 3: (a) Zero points of the vector $\phi^{r_{h}}$ shown in the $r_{h}-\tau$ plane with $a/r_{0}=1$, $Ar_{0}=1$ and $Pr_{0}^{2}=0.5$. There is one thermodynamically stable and one thermodynamically unstable four-dimensional Kerr-AdS C-metric black hole branch for any value of $\tau$. Obviously, the topological number is: $W=0$; (b) The arrows represent the unit vector field $n$ on a portion of the $r_{h}-\Theta$ plane for the four-dimensional Kerr-AdS C-metric black hole with $a/r_{0}=1$, $Ar_{0}=1$, $Pr_{0}^{2}=0.5$ and $\tau/r_{0}=20$. The zero points (ZPs) marked with black dots are at $(r_{h}/r_{0},\Theta)=(0.33,\pi/2)$, $(2.23,\pi/2)$, for ZP1 and ZP2, respectively. The blue contours $C_{i}$ are closed loops surrounding the zero points.
@@
Furthermore, in light of another intriguing finding presented in Ref. PRD108-084041 , which demonstrate that the topological number of an non-rotating, accelerating AdS black hole is consistently lower by $-1$ compared to its non-accelerating counterpart, a broader conjecture within the framework of general relativity is proposed: the topological number of an accelerating AdS black hole consistently differs by minus one from that of its corresponding non-accelerating AdS black hole. This conjecture will be further explored in section IV.2, where we calculate the topological number of the KN-AdS C-metric black hole and compare it to that of the KN-AdS black hole, thereby affirming the validity and relevance of our proposal.
@@
IV Topological classes of rotating charged accelerating black hole
@@
IV.1 KN C-metric black hole
@@
$Q=\frac{q}{K}\,,~{}\qquad\Phi=\frac{qr_{h}}{\alpha(r_{h}^{2}+a^{2})}\,,~{}\qquad\Xi=1+A^{2}(a^{2}+q^{2})\,,$
where $Q$ and $\Phi$ denote the electric charge and its corresponding electrostatic potential at the event horizon, respectively. Nevertheless, we can verify that both the differential and integral mass formulas are completely satisfied
@@
Next, we will discuss the topological number of the KN C-metric black hole. It is simple to derive the generalized off-shell Helmholtz free energy as
@@
Figure 4: (a) Zero points of the vector $\phi^{r_{h}}$ shown in the $r_{h}-\tau$ plane with $a/r_{0}=1$, $q/r_{0}=1$ and $Ar_{0}=0.5$. There is always one thermodynamically stable four-dimensional KN C-metric black hole for any value of $\tau$. Obviously, the topological number is: $W=1$; (b) The arrows represent the unit vector field $n$ on a portion of the $r_{h}-\Theta$ plane for the four-dimensional KN C-metric black hole with $a/r_{0}=1$, $q/r_{0}=1$, $Ar_{0}=0.5$ and $\tau/r_{0}=20$. The zero point (ZP) marked with black dot is at $(r_{h}/r_{0},\Theta)=(3.56,\pi/2)$. The blue contours $C$ is closed loop enclosing the zero point.
Considering $a/r_{0}=1$, $q/r_{0}=1$ and $Ar_{0}=0.5$ for the KN C-metric black hole, we plot the zero points of $\phi^{r_{h}}$ in the $r_{h}-\tau$ plane in Fig. 4, and the unit vector field $n$ on a portion of the $\Theta-r_{h}$ plane with $\tau/r_{0}=20$ in Fig. 4. Evidently, for any value of $\tau$, there exists a unique thermodynamically stable KN C-metric black hole. As shown in Fig. 4, the zero point is located at $(r_{h}/r_{0},\Theta)=(3.56,\pi/2)$. By analyzing the local properties of this zero point, we can readily determine the topological number $W=1$ for the KN C-metric black hole. Comparing this result with the corresponding finding for the Kerr C-metric black hole in section III.1, where $W=1$ as well, it becomes evident that the presence of the electric charge parameter does not influence the topological number of rotating, accelerating black holes. This observation further supports our initial conjecture, as proposed in Ref. PRD107-024024 , that the topological number remains unaffected by the charge in rotating black holes–a conclusion that not only extends to the case of rotating black holes in gauged supergravity theories PLB856-138919 , but also to the case of rotating, accelerating black holes. In addition, by comparing with the corresponding result for the KN black hole ($W=0$), we further validate the conjecture proposed in section III.1: the difference in topological numbers between any asymptotically flat accelerating black hole and its non-accelerating counterpart consistently equals one. This conclusion also holds true in the case of rotating, accelerating charged black holes.
IV.2 KN-AdS C-metric black hole
In this subsection, we consider the general KN-AdS C-metric black hole case. The metric, the Abelian gauge potential, and the metric functions are already given in Eqs. (6)-(II). The thermodynamic quantities of the KN-AdS C-metric black hole are JHEP0419096
@@ IV.2 KN-AdS C-metric black hole
Figure 5: (a) Zero points of the vector $\phi^{r_{h}}$ shown in the $r_{h}-\tau$ plane with $a/r_{0}=1$, $q/r_{0}=1$, $Ar_{0}=1$ and $Pr_{0}^{2}=0.5$. There is one thermodynamically stable and one thermodynamically unstable four-dimensional KN-AdS C-metric black hole branch for any value of $\tau$. Obviously, the topological number is: $W=0$; (b) The arrows represent the unit vector field $n$ on a portion of the $r_{h}-\Theta$ plane for the four-dimensional KN-AdS C-metric black hole with $a/r_{0}=1$, $q/r_{0}=1$, $Ar_{0}=1$, $Pr_{0}^{2}=0.5$ and $\tau/r_{0}=20$. The zero points (ZPs) marked with black dots are at $(r_{h}/r_{0},\Theta)=(0.42,\pi/2)$, $(2.17,\pi/2)$, for ZP1 and ZP2, respectively. The blue contours $C_{i}$ are closed loops surrounding the zero points.
In the following, we will investigate the topological number of the KN-AdS C-metric black hole. The generalized off-shell Helmholtz free energy is given by
@@
$\phi^{r_{h}}$ $=$ $\frac{\sqrt{1+A^{2}(a^{2}+q^{2})}}{6K[1-8\pi Pa^{2}/3+A^{2}(a^{2}+q^{2})]}\sqrt{1+\frac{A^{2}[-3+8\pi Pa^{2}-3A^{2}(a^{2}+q^{2})]}{8\pi P}}\Bigg{\{}3-\frac{3q^{2}}{r_{h}^{2}}-\frac{8\pi Pr_{h}^{2}(A^{2}r_{h}^{2}-3)}{(A^{2}r_{h}^{2}-1)^{2}}$
$+a^{2}\Bigg{[}4\pi P\bigg{(}\frac{1}{(Ar_{h}-1)^{2}}+\frac{1}{(Ar_{h}+1)^{2}}\bigg{)}\Bigg{]}\Bigg{\}}-\frac{2\pi A^{2}r_{h}(r_{h}^{2}+a^{2})}{K\tau(A^{2}r_{h}^{2}-1)^{2}}-\frac{2\pi r_{h}}{K\tau(1-A^{2}r_{h}^{2})}\,,$
@@
$\tau$ $=$ $-\frac{8\sqrt{2P}\pi^{\frac{3}{2}}r_{h}^{3}(1+A^{2}a^{2})\big{[}3-8\pi Pa^{2}+3A^{2}(a^{2}+q^{2})\big{]}}{\sqrt{\big{[}1+A^{2}(a^{2}+q^{2})\big{]}\big{[}8\pi P+A^{2}(8\pi Pa^{2}-3)-3A^{4}(a^{2}+q^{2})\big{]}}}\Big{\{}-3(r_{h}^{2}+8\pi Pr_{h}^{4})+3q^{2}(A^{2}r_{h}^{2}-1)^{2}$ (42)
$+A^{2}r_{h}^{4}\big{[}r_{h}^{2}(8\pi P-3A^{2})+6\big{]}+a^{2}\big{[}3(A^{2}r_{h}^{2}-1)^{2}-8\pi Pr_{h}^{2}(1+A^{2}r_{h}^{2})\big{]}\Big{\}}^{-1}$
@@
In Figs. 5 and 5, taking $a/r_{0}=1$, $q/r_{0}=1$, $Ar_{0}=1$, $Pr_{0}^{2}=0.5$ for the KN-AdS C-metric black hole, we plot the zero points of $\phi^{r_{h}}$ in the $r_{h}-\tau$ plane and the unit vector field $n$ with $\tau=20r_{0}$, respectively. Note that for these values of $a/r_{0}$, $q/r_{0}$, $Ar_{0}$ and $Pr_{0}^{2}$, there are one thermodynamically stable and one thermodynamically unstable KN-AdS C-metric black hole for any value of $\tau$. In Fig. 5, one can observe that two zero points are located at $(r_{h}/r_{0},\Theta)=(0.42,\pi/2)$ and $(2.17,\pi/2)$, respectively. As a result, the topological number $W=0$ for the KN-AdS C-metric black hole can be explicitly established in Figs. 5 and 5 via the local property of the zero point, which is different from that of the four-dimensional KN-AdS black hole ($W=1$) PRD107-084002 . Therefore, we further verify that the conjecture proposed in section III.2–”the topological number of an asymptotically AdS, accelerating black hole consistently differs by minus one from that of its corresponding asymptotically AdS, non-accelerating black hole”–is also applicable to the case of rotating, accelerating charged AdS black holes. Furthermore, since the KN-AdS C-metric black hole retains the same topological number as the KN black hole ($W=0$), we show that although both the acceleration parameter and the negative cosmological constant individually increase the topological number by one, their combined effect in the rotating charged black hole solution counteracts this increase.
V Concluding remarks